Compute the 2019 Federal income tax liability and the marginal and effective tax rates in each of the following independent cases.

Click here to access the 2019 tax rate schedule. If required, round the tax liability the nearest dollar. When required, round the average rates to four decimal places before converting to a percentage (i.e. .67073 would be rounded to .6707 and entered as 67.07%).

a. Chandler is single and reports taxable income of $115,400.

Tax liability: $
Marginal rate: %
Average rate: %

b. Lazare, a head of household, records taxable income of $53,000.

Tax liability: $
Marginal rate: %
Average rate: %
Answer #1
a
Tax liability: 21870 =14382+(115400-84200)*24%
Marginal rate: 24%
Average rate: 18.95% =21870/115400
b
Tax liability: 6098 =6065+(53000-52850)*22%
Marginal rate: 22%
Average rate: 11.51% =6098/53000